# EasyOcc: 3D Pseudo-Label Supervision for Fully Self-Supervised Semantic Occupancy Prediction Models

### v3

@@ III-A Generation of 3D Pseudo-Labels
This section presents a key contribution of this paper: the generation of 3D pseudo-labels from semantic maps of Grounded-SAM and depth maps from Metric3Dv2. This method enables loss computation directly in 3D voxel space, eliminating the need for view synthesis and aligning our approach with supervised training pipelines. The process is illustrated in Figure 2 and divided into three steps: semantic point cloud generation (Subsubsection III-A1), densification (Subsubsection III-A2), and voxelization (Subsubsection III-A3).
@@ III-A1 Semantic Point Cloud Generation
This stage will detail the generation of a semantic point cloud for an arbitrary sample. Semantic maps are sourced from the OccNeRF repository, which are generated using Grounded-SAM, while depth maps are generated with the Giant variant of Metric3Dv2 for optimal training performance [metric3dv2]. For each camera, $i\in\{1,\dots,6\}$, in a sample, given the corresponding semantic map $S_{i}\in\mathbb{R}^{H\times W}$, depth map $D_{i}\in\mathbb{R}^{H\times W}$, camera intrinsic matrix $K_{i}\in\mathbb{R}^{3\times 3}$, and camera-to-global transformation $\mathbf{T}_{\text{camera, }i}^{\text{global}}\in\mathbb{R}^{4\times 4}$, each arbitraty pixel $(u,v)\in[0,W)\times[0,H)$ is projected into the dehomogenised 3D global coordinates in Equation (1):
@@
To improve spatial accuracy, outlier removal is performed on $\mathcal{P}^{\text{sem}}$, yielding a consolidated semantic point cloud in the global coordinate frame. Outlier removal is implemented in the Open3D library [open3d], and is commonly used for preprocessing in 3D data analysis.
@@ III-A3 Semantic Point Cloud Voxelization
Once $\mathcal{P}_{T}^{\text{sem}}$ is obtained, it is voxelized according to the voxel bounds defined by the Occ3D ground truth: $40\text{m},-40\text{m},-1\text{m},40\text{m},40\text{m},5.4\text{m}]$, using a voxel resolution of $0.4\text{m}^{3}$, expressed in the ego-frame coordinate system of the current sample [occ3d]. Given the high density of $\mathcal{P}_{T}^{\text{sem}}$ due to the aggregation of many temporal samples, a voxel is considered occupied only if it contains a minimum of ten points; otherwise, it is treated as empty. This threshold helps mitigate the influence of stray points that could otherwise result in erroneous voxelization. For voxels classified as occupied, the semantic label is assigned based on the majority class among the contained points. The final output is a 3D pseudo-label designed for use in self-supervised semantic occupancy prediction models, specifically adapted for the Occ3D dataset.
III-B 3D Pseudo-Label Quality
In this section, we compare our generated 3D pseudo-labels with the ground-truth labels from Occ3D, both quantitatively and qualitatively.
@@ III-B1 Quantitative Analysis
Here, we examine the effect that aggregating temporal samples and thresholding occupancy has on likeness between our labels and the ground-truth labels in regards to the mIoU metric.
In Figure 3, we compare 3D pseudo-labels generated using varying numbers of aggregated temporal samples against the Occ3D ground-truth labels. The result follows a logarithmic trend, indicating saturation, where aggregating more temporal samples provides diminishing returns in mIoU. The optimal number of temporal samples is found to be 13, at which point we achieve the highest mIoU score of 13.58. These findings show that incorporating temporal samples improves the similarity of the pseudo-labels to the Occ3D ground truth. However, as the number of temporal samples increases, fewer instances can fully utilize them, for example, the 13th sample has only 12 predecessors. The maximum number of temporal samples is capped at 13 due to memory constraints.
Figure 3: Temporal Sample Aggregation Analysis: 3D pseudo-labels directly compared to ground-truth labels for various numbers of aggregated samples. A camera mask is applied to align with the model evaluation pipeline.
In Figure 4, we compare 3D pseudo-labels generated using different occupancy threshold values, ranging from 1 to 25. The highest performance is observed at a threshold of 3, indicating that even noisy points contribute useful information. For our experiments, we selected a threshold of 10 to balance slightly faster generation time with comparable accuracy.
Figure 4: Occupancy Threshold Analysis: 3D pseudo-labels compared to ground-truth labels for various threshold values in the generation processes. A camera mask is applied.
Figure 5: Occ3D Ground Truth (top) and Our 3D Pseudo-Labels(bottom): Visual comparison of four samples from the Occ3D dataset with their corresponding 3D pseudo-labels generated by our method.
@@ III-B2 Qualitative Analysis
In Figure 5, we compare four 3D pseudo-label training samples with their corresponding ground-truth labels. The pseudo-labels closely match the ground truth, accurately identifying key scene elements such as roads, vegetation, and buildings. Despite mitigation efforts such as outlier removal and occupancy thresholding, our pseudo-labels remain cluttered, primarily due to noise in the depth maps. However, as will be discussed in Subsection V-F, the model’s predicted outputs often appear smoother and more continuous than the ground-truth labels. In the rightmost sample, scene densification is lacking due to the sample being early in the sequence, resulting in limited aggregation of temporal data.
@@ IV Methodology II: Models
This section starts with Subsection IV-A, which outlines the loss function used for our 3D pseudo-labels, pseudo-loss. Subsection IV-B details the modifications made to existing models for integrating this loss function, while Subsection IV-C introduces our model, EasyOcc, which solely utilizes this loss function. The section concludes with Subsection IV-D, which explains the camera mask and describes the purpose of evaluating the models with and without it.
IV-A Pseudo Loss
Now that we have described the 3D pseudo-label generation process, we shall outline the loss function for using the labels during training. The loss function, pseudo-loss, will be used in the chosen modified models and EasyOcc. This loss consists of two distinct terms, as shown in (5), where $\lambda$ is a constant initialized at the start of training (an ablation on the value of $\lambda$ is discussed in Table VI). The formulation is adapted from GaussianOcc, where it was initially used as an optional component for training with ground-truth labels [gaussianocc].
As shown in Equation (6), geometry loss is composed of three separate losses: geometric scale loss, semantic scale loss, and Lovász softmax loss [lovasz].
All losses, including cross-entropy, are standard in semantic occupancy prediction, as they effectively penalize misclassifications and support class re-weighting to address dataset imbalance.
IV-B Modifications to Existing Models
Incorporating pseudo-loss into the three selected architectures requires considering how the loss function interacts with existing losses and also specific implementation details, as outlined below.
As described in Section II, SelfOcc predicts semantic occupancy using a two-step process: binary occupancy prediction (occocc), followed by a semantic voxel prediction (semsem) which together produce the final output. This structure introduces two key considerations: (1) a voxel may be classified as occupied by occocc but empty by semsem, resulting in it being considered unoccupied, and (2) a voxel may be classified as unoccupied by occocc but occupied by semsem, leading to it remaining unoccupied.
In OccNeRF, pseudo-loss is implemented alongside additional loss components to form the final loss function, as shown in Equation (8). The pseudo-loss serves as a complement to the three existing losses in the original OccNeRF model: $\mathcal{L}_{\text{regularisation}}$, $\mathcal{L}_{\text{reprojection}}$, $\mathcal{L}_{\text{sem}}$. These losses regulate voxel occupancy stability, multi-frame photometric consistency, and 2D semantic loss, respectively.
Given the similarity between GaussianOcc and OccNeRF, our pseudo-loss implementation follows the same approach in GaussianOcc, with one key difference: we omit $\mathcal{L}_{\text{sem}}$ due to sporadic NaN gradients in the convolutional layers of the image encoder during training. The cause of this issue remains unknown. The resulting loss function is defined in Equation (9).
IV-C EasyOcc
Integrating pseudo-loss in EasyOcc is straightforward, as the framework relies exclusively on this signal for learning. The continuous and dense scene representation enables effective learning in conjunction with our 3D pseudo-labels.
The model is a simplified variant of GaussianOcc, with its architectural flow illustrated in Figure 1. Multi-view camera images are processed through a ResNet-101 image encoder to extract high-level features, which have shown robust performance across BEV models [simplebev, bevcar] and semantic occupancy prediction frameworks [occnerf, gaussianocc, daocc]. Parameter-free bilinear sampling projects these features into 3D space, which are then passed through a 3D CNN to enhance spatial reasoning and produce semantic logits.
Up to this point, the pipeline mirrors OccNeRF and GaussianOcc. However, our approach to loss computation diverges by directly computing loss against the generated 3D pseudo-labels, eliminating the need for depth estimation, multi-frame consistency, and view synthesis, thus reducing training complexity and duration.
IV-D Camera Mask
Comprehensive scene modeling requires an understanding of the full surrounding environment. In prior occupancy datasets [surroundocc, openocc], evaluation is conducted across the entire voxel grid, compelling models to accurately predict both occupancy and semantics for the entire scene.
Figure 6: Camera Mask Visualization: Semantic occupancy grid (left), and corresponding boolean camera mask (right). Camera mask includes empty voxels.
TABLE I: Model Configurations: Rendering time refers to the time required to render semantics, depth, or features for a single sample during training. Training time denotes the time per training epoch; the plus symbol (+) indicates the additional time incurred due to the inclusion of our pseudo-loss. * For OccNeRF, the parameter count includes the NeRF rendering module, which is used during training only. ** For GaussianOcc, the parameter count includes the pose estimation module, which is used during training only. (-) indicates the data was not listed in the paper or repository. $T$-OccCamera and AutoOcc are not listed as they require no training.
Model Backbone Model Parameters Image Size Epochs Rendering Time Training Time
@@ IV-D Camera Mask
[…]
In contrast, models assessed on the Occ3D benchmark are evaluated solely on voxels visible from the current camera view. These visible voxels, referred to as the camera mask, are determined by ray casting from each surround-view camera in the current sample to the voxel grid [occ3d]. This evaluation strategy limits overall scene comprehension and reduces penalties for issues such as overprediction and object duplication. Given the availability of temporal cues, the ability to infer occluded areas, and the use of priors, such as typical object dimensions, it is important to also evaluate models over the full voxel grid. Figure 6 illustrates a sample’s semantic occupancy grid alongside its boolean camera mask. While the entire scene contains 640,000 voxels, models evaluated under Occ3D are assessed on only 67,761 of these.
As will be demonstrated in Subsection V-D, our 3D pseudo-labels enhance performance not only under the standard camera mask evaluation but also significantly boost results when evaluated over the full voxel grid. While previous works primarily report performance using the camera mask, we extend this by providing full-voxel grid evaluations for all models, highlighting that our 3D pseudo-labels contribute to a more holistic understanding of the scene.
@@ V Results
In this section, we present the main results of this paper. We begin with a description of the dataset and evaluation metrics in Subsection V-A, and a brief explanation of model configurations in Subsection V-B. In Subsections V-C and V-D, we examine the effectiveness of the deployment of our 3D pseudo-labels, evaluated with and without the camera-mask, respectively. This is followed by an ablation study for the EasyOcc in Subsection V-E. Subsection V-F analyzes the qualitative results based on the main results table. The section concludes with a summary and discussion of all findings in Subsection V-G.
@@ V-A Dataset and Evaluation Metrics
All models are trained and evaluated on the Occ3D dataset, which consists of 600 training scenes and 150 validation scenes from the nuScenes dataset. The voxel space is bounded by $40\text{m},-40\text{m},-1\text{m},40\text{m},40\text{m},5.4\text{m}]$, with a voxel size of $0.4\text{m}^{3}$.
Models are evaluated using the Intersection over Union (IoU) and mean Intersection over Union (mIoU) metrics. IoU, defined in Equation (10), reflects the model’s ability to capture overall spatial structure through occupancy. The mIoU metric, shown in Equation (11), computes the average IoU across all semantic classes, excluding the empty class (class index 17).
@@
$\text{mIoU}=\frac{1}{C-1}\sum_{\begin{subarray}{c}c=0\\ c\neq 17\end{subarray}}^{C}\frac{TP_{c}}{TP_{c}+FP_{c}+FN_{c}}$ (11)
@@ V-B Model Configurations
In Table I, we compare all model configurations. Rendering time and depth estimation add overhead, with Gaussian Splatting being the most efficient due to its rasterization-based rendering [gaussian_splatting]. EasyOcc avoids these rendering methods during training, resulting in reduced training time. While incorporating pseudo-loss increases epoch training time, the impact varies by model, with OccNeRF experiencing the largest increase of (+31m) due to its overall slower training.
TABLE II: EasyOcc Inference Time: Time required for each component of the inference pipeline.
@@
TABLE III: State of the Art Comparison with the Camera Mask: FPS denotes frames per second, representing the time taken to process a single sample. IoU and mIoU refer to Intersection over Union and mean Intersection over Union, respectively. The best result in each category is highlighted in bold, while the second-best is underlined. * OccNeRF implements 2D semantic loss, whereas GaussianOcc does not.
@@
Model FPS IoU mIoU $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SelfOcc [selfocc] 7.4 44.05 9.66 0.20 0.52 6.71 10.42 0.00 0.10 2.12 0.00 0.00 7.72 56.13 26.92 25.68 13.38 4.62 OccNeRF [occnerf] 5.4 46.43 9.73 0.68 1.81 6.59 6.59 3.68 0.34 2.93 3.19 2.90 6.64 52.80 23.99 24.97 18.62 9.69 GaussianOcc [gaussianocc] 5.4 42.91 9.94 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 20.10 17.58 8.61 10.29 TT-OccCamera [ttgaussocc] 0.7 - 1184 0.00 5.90 8.94 12.58 2.75 9.67 4.71 4.04 0.00 8.77 55.65 26.49 30.20 15.13 16.57 GaussTR [gausstr] 0.3 44.54 12.27 6.50 8.54 21.77 24.27 6.26 15.48 7.94 1.86 6.10 17.16 36.98 17.21 7.16 21.18 9.99 GaussianFlowOcc [gaussianflowocc] 10.2 46.91 15.07 7.23 9.33 17.55 17.94 4.50 9.32 8.51 10.66 2.00 11.80 63.89 31.11 35.12 14.64 12.59 AutoOcc [autoocc] - 83.01 17.87 2.70 10.45 7.81 20.42 5.79 17.58 18.50 24.25 4.23 12.88 55.54 24.23 27.14 35.62 36.61 SelfOcc [selfocc] + Ours 7.4 34.50 13.43 1.66 5.43 14.51 22.22 2.60 6.39 15.38 8.90 1.00 12.77 54.97 26.81 21.55 11.34 9.29 OccNeRF [occnerf] + Ours * 5.4 38.46 14.09 1.85 8.18 16.66 22.12 1.01 7.74 14.74 12.84 0.98 13.76 55.91 27.96 22.73 15.77 17.20 GaussianOcc [gaussianocc] + Ours * 5.4 38.78 13.89 1.74 5.85 16.22 22.34 2.30 8.39 15.71 10.18 0.99 13.36 55.29 27.36 23.41 15.96 16.97 EasyOcc (Ours) RN-101 5.4 38.86 13.86 1.90 6.67 15.09 21.68 2.70 8.06 15.28 11.08 1.36 12.81 55.78 27.85 22.06 16.07 17.30
@@
In Table II, we report the inference time (in milliseconds) for each component of the EasyOcc model. Grid Sampling refers to downsampling the contracted coordinate voxel grid to align with the dimensions of the Occ3D ground-truth labels, a technique introduced in OccNeRF [occnerf]. Preliminary experiments showed that retaining the contracted coordinate system, rather than modeling the scene in the Occ3D output space, improved performance.
@@
Self-supervised models that utilize LiDAR during training or inference are excluded from our comparison [ago, ttgaussocc, minkocc, veon] as our method is a camera-only pipeline, consistent with the models used for comparison in this study. Training and inference are performed on four NVIDIA A100-SXM4-40GB GPUs.
V-C Main Results: Camera Mask

### v4

@@ III-A GENERATION OF EASY3D-LABELS
This section presents the key contribution of this paper: the generation of our 3D pseudo-labels from semantic maps of Grounded-SAM [37] and depth maps from Metric3Dv2 [14]. This method enables loss computation directly in 3D voxel space, eliminating the need for view synthesis and aligning our approach with supervised training pipelines. The process is illustrated in Figure 1 and divided into three steps: semantic point cloud generation (Subsubsection III-A1), densification (Subsubsection III-A2), and voxelization (Subsubsection III-A3).
@@ III-A1 SEMANTIC POINT CLOUD GENERATION
This stage will detail the generation of a semantic point cloud for an arbitrary sample. Semantic maps are sourced from the OccNeRF repository [55], which are generated using Grounded-SAM, while depth maps are generated with the Giant variant of Metric3Dv2 for optimal training performance [14]. For each camera, $i\in\{1,\dots,6\}$, in a sample, given the corresponding semantic map $S_{i}\in\mathbb{R}^{H\times W}$, depth map $D_{i}\in\mathbb{R}^{H\times W}$, camera intrinsic matrix $K_{i}\in\mathbb{R}^{3\times 3}$, and camera-to-global transformation $\mathbf{T}_{\text{camera, }i}^{\text{global}}\in\mathbb{R}^{4\times 4}$, each arbitraty pixel $(u,v)\in[0,W)\times[0,H)$ is projected into the dehomogenised 3D global coordinates in Equation (1):
@@
To improve spatial accuracy, outlier removal is performed on $\mathcal{P}^{\text{sem}}$ using the Open3D library [62], yielding a consolidated semantic point cloud in the global coordinate frame.
@@ III-A3 SEMANTIC POINT CLOUD VOXELIZATION
Once $\mathcal{P}_{T}^{\text{sem}}$ is obtained, it is voxelized to obtain the final 3D pseudo-label. The bounds are defined by the Occ3D-nuScenes [45] ground truth: $40\text{m},-40\text{m},-1\text{m},40\text{m},40\text{m},5.4\text{m}]$, using a voxel resolution of $0.4\text{m}^{3}$, expressed in the ego-frame coordinate system of the current sample. Given the high density of $\mathcal{P}_{T}^{\text{sem}}$ due to the aggregation of many temporal samples, a voxel is considered occupied only if it contains a minimum of ten points; otherwise, it is treated as empty. This threshold helps mitigate the influence of stray points that could otherwise result in erroneous voxelization. For voxels classified as occupied, the semantic label is assigned based on the majority class among the contained points.
III-B EASY3D-LABELS QUALITY
In this section, we compare our Easy3D-Labels with the ground-truth labels from Occ3D-nuScenes [6], both quantitatively and qualitatively.
@@ III-B1 QUANTITATIVE ANALYSIS
In Figure 2, we compare Easy3D-Labels using varying numbers of aggregated temporal samples against the Occ3D-nuScenes ground-truth labels. The result follows a logarithmic trend, indicating saturation, where aggregating more temporal samples provides diminishing returns in mIoU. The optimal number of temporal samples is found to be 13, at which point we achieve the highest mIoU score of 15.4. These findings show that incorporating temporal samples improves the similarity of the pseudo-labels to the Occ3D ground truth. The maximum number of temporal samples is capped at 13 due to memory constraints.
Figure 2: Temporal sample aggregation: Easy3D-Labels compared to Occ3D-nuScenes [45] labels for various numbers of aggregated samples.
In Figure 3, we compare our labels generated using different occupancy threshold values, ranging from 1 to 25. The highest performance is observed at a threshold of 3, indicating that even noisy points contribute useful information. For our experiments, we selected a threshold of 10 to balance slightly faster generation time with comparable accuracy.
Figure 3: Occupancy threshold: Easy3D-Labels compared to Occ3D-nuScenes [45] labels for various threshold values in the generation processes.
@@ III-B2 QUALITATIVE ANALYSIS
In Figure 4, we compare four Easy3D-Labels training samples with their corresponding ground-truth labels. The pseudo-labels closely match the ground truth, accurately identifying key scene elements such as roads, vegetation, and buildings. Despite mitigation efforts such as outlier removal and occupancy thresholding, there exist incorrectly labelled voxels primarily due to noise in the depth maps. However, as will be discussed in Subsection V-F, the model’s predicted outputs often appear smoother and more continuous than the ground-truth labels. Densification is lacking in the rightmost sample due to the sample being early in the sequence, resulting in limited aggregation of temporal data.
Figure 4: Pseudo-label comparison: Occ3D-nuScenes [45] ground truth (top) and our Easy3D-Labels (bottom)
@@ IV METHODOLOGY II: MODELS
This section introduces the pseudo-loss, $\mathcal{L}_{\text{Pseudo}}$, in Subsection IV-A, model modifications in Subsection IV-B, and our model, EasyOcc, in Subsection IV-C.
TABLE I: Model configurations: Rendering time denotes the time to render semantics, depth, or features per sample during training, while training time is reported per epoch; (+) indicates additional time from pseudo-loss. * OccNeRF parameters include the NeRF rendering module (training only). ** GaussianOcc parameters include the pose estimation module (training only).
Method Backbone Model Parameters Image Size Epochs Rendering Time Training Time
@@
IV-A PSEUDO LOSS
Model supervision using Easy3D-Labels will be facilitated by the pseudo-loss function, $\mathcal{L}_{\text{Pseudo}}$. This loss consists of two distinct terms, as shown in (5), where $\lambda$ is a constant initialized at the start of training (an ablation on the value of $\lambda$ is discussed in Table VII). The formulation is adapted from GaussianOcc, where it was initially used as an optional component for training with ground-truth labels [9].
As shown in Equation (6), geometry loss is composed of three separate losses: geometric scale loss, semantic scale loss, and Lovász softmax loss [3].
All losses, including cross-entropy, are standard in semantic occupancy estimation, as they effectively penalize misclassifications and support class re-weighting to address dataset imbalance.
IV-B MODIFICATIONS TO EXISTING MODELS
Incorporating $\mathcal{L}_{\text{Pseudo}}$ into the three selected architectures requires considering how the loss function interacts with existing losses and also specific implementation details, as outlined below.
SelfOcc: As described in Section II, SelfOcc predicts semantic occupancy using a two-step process: binary occupancy prediction (occocc), followed by a semantic voxel prediction (semsem), which together produce the final output. This structure introduces two key considerations: (1) a voxel may be classified as occupied by occocc but empty by semsem, resulting in it being considered unoccupied, and (2) a voxel may be classified as unoccupied by occocc but occupied by semsem, leading to it remaining unoccupied.
OccNeRF: Here, we implement pseudo-loss alongside additional loss components to form the final loss function, as shown in Equation (8). The pseudo-loss serves as a complement to the three existing losses in the original OccNeRF model: $\mathcal{L}_{\text{regularisation}}$, $\mathcal{L}_{\text{reprojection}}$, $\mathcal{L}_{\text{sem}}$. These losses regulate voxel occupancy stability, multi-frame photometric consistency, and 2D semantic loss, respectively.
GaussianOcc: Given the similarity between GaussianOcc and OccNeRF, our pseudo-loss implementation follows the same approach in GaussianOcc, with one key difference: we omit $\mathcal{L}_{\text{sem}}$ due to sporadic NaN gradients in the convolutional layers of the image encoder during training. The cause of this issue remains unknown. The resulting loss function is defined in Equation (9).
Figure 5: EasyOcc model architecture: Image features are extracted and then processed in voxel space by 3D convolutions prior to pseudo-loss computation against our 3D pseudo-labels, Easy3D-Labels.
IV-C EASYOCC
Integrating $\mathcal{L}_{\text{Pseudo}}$ in EasyOcc is straightforward, as the framework relies exclusively on this signal for learning. The continuous and dense scene representation enables effective learning in conjunction with our 3D pseudo-labels.
The model is a simplified variant of GaussianOcc, with its architectural flow illustrated in Figure 5. Multi-view camera images are processed through a ResNet-101 image encoder to extract high-level features, which have shown robust performance across BEV models [11] and semantic occupancy prediction frameworks [9]. Parameter-free bilinear sampling projects these features into 3D space, which are then passed through a 3D CNN to enhance spatial reasoning and produce semantic logits. EasyOcc eliminates the need for depth estimation, multi-frame consistency,and novel view synthesis, thus reducing training complexity and duration.
@@ V RESULTS
In this section, we present the main results. Subsections V-A and V-B describe the dataset, metrics, and configurations, followed by evaluations of Easy3D-Labels for mIoU and RayIoU in Subsections V-C and V-D. We then provide an ablation study of EasyOcc in Subsection V-E and qualitative analysis in Subsection V-F.
@@ V-A DATASET AND EVALUATION METRICS
We evaluate all models on mIoU and RayIoU on the Occ3D-nuScenes dataset [45], and on RayIoU only on the OpenOccv2 dataset [46], as it does not contain a camera mask, which permits fair mIoU evaluation for self-supervised models. However, OpenOccv2 provides denser labels, which leads to a fairer evaluation on the RayIoU metric. Both datasets consist of 600 training scenes and 150 validation scenes from the nuScenes dataset. The voxel space is bounded by $40\text{m},-40\text{m},-1\text{m},40\text{m},40\text{m},5.4\text{m}]$, with a voxel size of $0.4\text{m}^{3}$.
For evaluation metrics, we utilize Intersection over Union (IoU), mean Intersection over Union (mIoU), and RayIoU. IoU, defined in Equation (10), reflects the model’s ability to capture overall spatial structure through occupancy. The mIoU metric, shown in Equation (11), computes the average IoU across all semantic classes, excluding the empty class.
@@
$\text{mIoU}=\frac{1}{C}\sum_{\begin{subarray}{c}c=1\end{subarray}}^{C}\frac{TP_{c}}{TP_{c}+FP_{c}+FN_{c}}$ (11)
@@
RayIoU is defined similarly to mIoU, but instead of being a voxel-wise metric, it is a ray metric. Introduced in SparseOcc [44], RayIoU aims to resolve the harsh penalisation of incorrect depth estimations and the overcompensation of overprediction of voxels for inflating the mIoU score. A ray is cast from the LiDAR sensor position in both the ground-truth and predicted voxel gird and it is labelled correct if both ray depths are within a threshold and are the same class. The equation is:
$\text{RayIoU}=\frac{1}{C}\sum_{\begin{subarray}{c}c=1\end{subarray}}^{C}\frac{TP_{c}}{TP_{c}+FP_{c}+FN_{c}}$ (12)
@@ V-B MODEL CONFIGURATIONS
In Table I, we compare all model configurations. Rendering time and depth estimation add overhead, with Gaussian Splatting being the most efficient due to its rasterization-based rendering [21]. EasyOcc avoids these rendering methods during training, resulting in reduced training time. While incorporating pseudo-loss increases epoch training time, the impact varies by model, with OccNeRF experiencing the largest increase of (+31m) due to its overall slower training.
TABLE II: EasyOcc inference time breakdown.
@@
TABLE III: State-of-the-art comparison on the Occ3D-nuScenes [45] dataset: FPS denotes frames per second, indicating processing time per sample. IoU and mIoU refer to Intersection over Union and mean Intersection over Union. Grey rows indicate SOTA methods for reference. The best result for each model (compared to its variant with our labels) is highlighted in bold. * OccNeRF uses 2D semantic loss, while GaussianOcc does not.
@@
Method FPS IoU mIoU $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation DistillNeRF [48] 2.8 29.1 10.1 1.4 2.1 10.2 10.1 2.6 2.0 5.5 4.6 1.4 7.9 43.0 16.9 15.0 14.1 15.1 GaussTR [19] 0.3 44.5 13.8 6.5 8.5 21.8 24.3 6.3 15.5 7.9 1.9 6.1 17.2 37.0 17.2 7.2 21.2 10.0 TT-Occ [57] 0.7 - 16.7 21.5 10.5 10.7 14.7 11.9 12.3 9.7 12.2 4.4 7.9 48.3 23.7 28.3 14.1 20.2 GaussianFlowOcc [5] 10.2 46.9 17.1 6.8 9.7 19.0 17.2 4.2 11.8 9.3 10.3 1.8 12.3 61.0 31.2 34.8 14.7 12.4 AutoOcc [63] - 83.0 20.9 12.7 10.5 7.8 20.4 5.8 17.6 18.5 24.3 4.2 12.9 55.5 24.2 27.1 35.6 36.6 SelfOcc [17] 7.4 44.1 10.3 0.2 0.5 6.7 10.4 0.0 0.1 2.1 0.0 0.0 7.7 56.1 26.9 25.7 13.4 4.6 + Ours 7.4 34.5 14.4 1.7 5.4 14.5 22.2 2.6 6.4 15.4 8.9 1.0 12.8 55.0 26.8 21.6 11.3 9.3 OccNeRF [55] 5.4 46.4 11.0 0.7 1.8 6.6 6.6 3.7 0.3 2.9 3.2 2.9 6.6 52.8 24.0 25.0 18.6 9.7 + Ours 5.4 38.5 16.0 1.9 8.2 16.7 22.1 1.0 7.7 14.7 12.8 1.0 13.8 55.9 28.0 22.7 15.8 17.2 GaussianOcc [9] 5.4 42.9 11.3 1.8 5.8 14.6 13.6 1.3 2.8 8.0 9.8 0.6 9.6 44.6 20.1 17.6 8.6 10.3 + Ours 5.4 38.8 15.7 1.7 5.9 16.2 22.3 2.3 8.4 15.7 10.2 1.0 13.4 55.3 27.4 23.4 16.0 17.0 EasyOcc (Ours) 5.4 38.9 15.7 1.9 6.7 15.1 21.7 2.7 8.1 15.3 11.1 1.4 12.8 55.8 27.9 22.1 16.1 17.3
@@
In Table II, we report the inference time (in milliseconds) for each component of the EasyOcc model. Grid Sampling refers to downsampling the contracted coordinate voxel grid to align with the dimensions of the Occ3D ground-truth labels, a technique introduced in OccNeRF [55]. Preliminary experiments showed that retaining the contracted coordinate system, rather than modeling the scene in the Occ3D output space, improved performance.
@@
Self-supervised models that utilize LiDAR during training or inference are excluded from our comparison [23, 57, 42, 61] as our method is a camera-only pipeline, consistent with the models used for comparison in this study. Training and inference are performed on four NVIDIA A100-SXM4-40GB GPUs.
V-C MAIN RESULTS: MIOU